OpenStudy (anonymous):

(y ^ 2-xy) dx+x ^ 2dY = 0 y(1)=1

OpenStudy (anonymous):

Make it exact to find a general solution.

OpenStudy (anonymous):

resolver a equação homogenea

OpenStudy (anonymous):

\[x ^{2}\frac{ dy }{ dx }=-y ^{2}+xy,\frac{ dy }{ dx }=\frac{ -y ^{2}+xy }{x ^{2}}\] put y=vx \[\frac{ dy }{ dx }=v+x \frac{ dv }{ dx }\] \[v+x \frac{ dv }{ dx }=\frac{- v ^{2}x ^{2}+vx ^{2} }{x ^{2} }\] \[=-v ^{2}+v\] \[x \frac{ dv }{dx }=-v ^{2}+v-v=-v ^{2}\] separate the variables and integrate.